# Dredge Word, Social Media, and Webgraph Networks for Unreliable Website Classification and Identification

### v1

@@ Abstract
In an attempt to mimic the complex paths through which unreliable content spreads between search engines and social media, we explore the impact of incorporating both webgraph and large-scale social media contexts into website credibility classification and discovery systems. We further explore the usage of what we define as dredge words on social media—terms or phrases for which unreliable domains rank highly. Through comprehensive graph neural network ablations, we demonstrate that curriculum-based heterogeneous graph models that leverage context from both webgraphs and social media data outperform homogeneous and single-mode approaches. We further demonstrate that the incorporation of dredge words into our model strongly associates unreliable websites with social media and online commerce platforms. Finally, we show our heterogeneous model greatly outperforms competing systems in the top-k identification of unlabeled unreliable websites. We demonstrate the strong unreliability signals present in the diverse paths that users follow to uncover unreliable content, and we release a novel dataset of dredge words.
Dredge Word, Social Media, and Webgraph Networks for Unreliable Website Classification and Identification
Evan M. Williams  and Peter Carragher  and Kathleen M. Carley Carnegie Mellon University 5000 Forbes Ave. USA, Pittsburgh, PA, 15213-3890 {emwillia, pcarragh, carley}@andrew.cmu.edu
1 Introduction
@@
On February 24, 2022, the day Russia began its invasion of Ukraine, Jacob Creech, a fringe Qanon conspiracy theorist, posted a series of unsubstantiated claims on Twitter. These tweets insinuated that the US had created COVID-19, and implied Russia’s invasion of Ukraine was actually being carried out to shut down US-funded “biolabs” in Ukraine and to prevent another global pandemic League (2022). Hours after the tweet, the conspiratorial website InfoWars published an article promoting the tweet, crediting its author for uncovering an “ulterior motive theory” League (2022). The tweet began circulating to other platforms as Google queries for “US biolabs” and “Ukraine Biolabs” spiked in the following days  [^1] https://trends.google.com/trends/explore?date=today%205-y&geo=US&q=Ukraine%20biolabs&hl=en. As only conspiratorial sites covered the tweet in its early hours, those would have likely been the only relevant search results until Snopes debunked the claim later that day Evon (2022).
@@
In some ways, this is a success story of the current paradigm of combatting misinformation; fact-checkers responded rapidly to a quickly-spreading falsehood. In other ways, this highlights the challenges present in reactive misinformation interventions. Despite quick debunking by fact-checkers, the conspiracy was nonetheless echoed by established news sources like Fox News, and later amplified on the floor of the US SenateLeague (2022). Researchers later found that a coordinated campaign of social media accounts boosted the narrative on Twitter Alieva et al. (2022). As of 2024, despite the underlying claims having been debunked by many fact-checkers Chappell and Yousef (2022); Parachini (2022), a search engine audit study found that Google, Bing, and Yandex all returned URLs on their first page of results that promote the “Ukraine Biolabs” conspiracy Kuznetsova et al. (2024). This story also highlights an important and understudied dimension of misinformation spread: the interaction between social media and search engines.
@@
An alternative to reactive fact-checking approaches are proactive algorithmic content moderation approaches. These can include modifying the recommendation and ranking systems to decrease the reach and virality of unreliable information sources. In search engines, this can mean downranking articles from unreliable domains, and on social media, this can mean ranking posts containing unreliable domains lower in user newsfeeds. For proactive approaches to succeed, platforms need systems that can rapidly detect and discover unreliable domains. Both classification and detection are important as unreliable websites can, and do, deploy a number of tactics to avoid blacklists Carragher et al. (2024). In this work, we present a novel approach to the tasks of unreliable domain detection and discovery that leverages signals from both large-scale social media data and webgraph data.
@@
Directing users to search engines can be an effective tactic to spread misinformation as, since 2016, global surveys have consistently found that individuals report higher trust in search engines than traditional media Barometer (2024); McDuling (2015). Thus by pointing users to search engines, either explicitly or implicitly, bad actors can create a false sense of content reliability. Little is understood about these pathways, and research has only recently begun exploring the ways social media users end up in data voids. The concepts of problematic queries Golebiewski and Boyd (2019) and keyword signaling Tripodi (2019) are highly related, but have only been explored in case studies. The recently-proposed concept of search directives provides a clear and observable path that could direct users to unreliable content on search engines. A search directive is content that is explicitly intended to prompt an online search, e.g., user $a$ directs user $b$ to "look up Chemtrails on Google" Robertson et al. (2023). In the context of data voids, search directives as a concept have two limitations: First, while search directives can direct users to unreliable content, they can also direct users to innocuous or helpful information. Second, as the search trends for “Ukraine Biolabs” demonstrate, users can be driven to unreliable content through search engines by querying phrases that appear without explicit search directives. To help fill these gaps, we propose the concept of dredge words—terms or keyphrases for which unreliable domains rank highly on search engines.
By attempting to explicitly incorporate each of the paths that users take to unreliable websites into GNNs, we seek to explore the dynamics that connect the spread of unreliable content on social media and search engines. First, we attempt to demonstrate domain (un)reliability signal present in large-scale social media data and small-scale webgraph data with varying levels of context. With respect to website credibility classification, we use graph neural networks to demonstrate the domain reliability signal in user-domain interaction networks on large-scale twitter data, on small-scale webgraph networks, and homogeneous and heterogeneous combinations of the two. Through comprehensive graph neural network ablations, we demonstrate that heterogeneous graph models that leverage context from both webgraphs and social media data outperform homogeneous and single-mode approaches. We additionally demonstrate that designing a training curriculum based on relative website reliability scores further increases model performance.
@@
Finally, we provide the first exploration of dredge words on social media—terms or keyphrases for which unreliable domains rank highly in search engines. We demonstrate the usefulness of these dredge words in an unreliable domain classification and discovery process that mimics how social media users may transition from social platforms to the consumption of misinformation via search engines. Surprisingly, we also find that dredge words frequently surface social media URLs in top SERP positions, thus providing a bidirectional path that often leads back to social media. We show our heterogeneous model greatly outperforms competing systems in the top-k discovery of unlabeled unreliable websites. We publicly release the code, webgraph data collected for this project, a dataset of 3,939 dredge words for 46 unreliable news domains, and the resulting SERPs for each of the dredge word queries [^2] Artefacts will be publicly released following review period.
2 Related Works
2.1 Social Media
@@
Classifying individual texts or articles is a core problem addressed by research in misinformation detection. Such detection methods have typically relied on website content and social media data Castelo et al. (2019); Chen and Freire (2020); Silva et al. (2021). Castelo et al. proposed a topic-agnostic detection system Castelo et al. (2019) that identifies unreliable articles based on Linguistic Inquiry and Word Count features. Chen and Freire adopted the method for the task of unreliable domain discovery Chen and Freire (2020). Their discovery system capitalized on user tendencies within the social graph, wherein a user who tweets a URL from a known unreliable source is likely to also tweet URLs from yet unknown sources. Similarly, Silva et al. combine website content and social context resulting in a misinformation page detection method that leverages heterogeneous data types Silva et al. (2021), with a focus on early detection across a broad range of topics.
2.2 Webgraphs
The predictive power of webgraphs and social media data have been demonstrated separately in several contexts. Aswani et al. detect SEO manipulation by link-building sites by clustering on Pagerank score and Domain Authority Aswani et al. (2021). For detecting news site bias, Aires et al. Patricia Aires et al. (2019) scrape cross-links from a list of prominent news sites from Media Bias Fact Check (Zandt, 2022). Sehgal et al. explore a case where misinformation was spread through coordinated hyperlink and social media networks Sehgal et al. (2021) on a dataset of 1.4k URLs. Another recent case study demonstrated the manipulated webgraph linkages of unreliable pseudo-thinktanks Williams and Carley (2023). More recently, Carragher et al. (2024) proposed a webgraph-based model for unreliable domain detection and discovery tasks. For discovery, the authors use backlinking domains in the webgraph in a snowball sampling approach. The authors demonstrate that GNN models trained on SEO attributes and webgraph data proved effective for detection.
2.3 Combining Webgraphs and Social Media Context
Zhang and Cabage show that social media and SEO promotion have different strengths; they find that social-sharing results in immediate but short-term boosts to traffic, while the benefits of link-building are slower to realize, but last longer Zhang and Cabage (2017). This hints that a combination of both webgraph and social media data should be used in investigating misinformation sources. However, to our knowledge, this is the first work that attempts to combine the two contexts for unreliable domain classification or discovery.
3 Data
We construct datasets based on the paths that users can take to unreliable websites. In our social media and webgraph data, we attempt to capture explicit paths users take to unreliable websites—i.e. clicking a URL. The dredge words we collect for a subset of the unreliable domains represent implicit paths.
3.1 URL Labels
We use the domain credibility labels published by Lin et al. Lin et al. (2023). The authors of the work align the domain reliability ratings of all domains ranked by 6 expert groups and run imputation followed by principal component analysis to generate aggregate ratings for 11,520 domains Lin et al. (2023). To our knowledge, this is the most comprehensive public domain reliability rating list as of spring of 2024. The domains in this list are ranked by the calculated first principal component where low scores correspond to high expert agreement on unreliability and high scores correspond to expert agreement on reliability. While these scores are very sensible across wide principal component score gaps, locally the relative orderings are less clear. Treating website reliability as continuous also muddies the interpretability and the discovery process, as there’s not a transparent reason a website should have a score of 0.571 as opposed to 0.570. To address this issue, we elected to binarize the website reliability labels. After manual inspection of the data, we elected to consider the bottom two quintiles as unreliable (corresponding to a principal component threshold of 0.5162).
3.2 Webgraph and features
As extracting complete webgraph information for these domains is computationally challenging, researchers have generally needed to use third-party tools to extract local webgraph attributes. We use the SEO toolkit service Ahrefs333ahrefs.com to extract the 10 domains which link to each target domain at the highest volume (the highest-volume back-linking domains). We chose Ahrefs because researchers have previously shown the usefulness of its data in domain reliability classification and discovery applications Carragher et al. (2024); Williams and Carley (2023). We were unable to pull backlinks for 193 of the domains. A random sample of 10 of 193 domains and found 9 of them were either dead or unreachable. We therefore elected to drop these 193 domains, leaving us with 11,327 unique reliability-labeled domains and 32,431 unique unlabeled backlinking domains for a total of 43,758 domains. For each of these 43,758 domains, we pull 23 attributes, which for an individual domain, contains fields like the total number of backlinks and outlinks, number of backlinks coming from .edu or .gov domains, and number of referring pages  [^4] https://ahrefs.com/api/documentation/metrics.
Similar to the backlink network constructed in Carragher et al. (2024), there is clear assortativity across label-reliability groups as can be seen in Figure 1. The prevalence of unlabeled nodes in our network overwhelms a node attribute assortativity coefficient calculation $r_{A}=-0.336$. We re-calculate node attribute assortativity on an induced subgraph that only maintains edges between labeled nodes. The subgraph contains less structure and 176 components, but nevertheless displays strong positive associativity $r_{A}=0.376$. In other words, domains of the same (binary) reliability are more likely to link to one another than linking to a domain of a different reliability. This provides a strong justification for the use of network-based models.
3.3 Twitter Data
@@
For social media context, we use a Twitter dataset constructed by querying COVID-related keywords555coronavirus, Wuhan virus, Wuhanvirus, 2019nCoV, NCoV, NCoV2019, covid-19, covid19, covid 19 via Twitter’s streaming API between January 29, 2020 and June 26, 2022. Due to server issues and API limitations, 121 days over the time period have partial or missing data. However, these gaps are spread relatively evenly over the time period, and still provide strong coverage. Our final Twitter dataset contained 3.6 billion extracted tweets.
@@
From this set of 3.6 billion tweets, we pull all Tweets that mention or link to one of our 11,520 websites, resulting in a dataset of 320M tweets that link to over 840k unique domains. Of the original 11,520 labeled domains, only 5,504 appeared within the Twitter data. To reduce the noise in the dataset we undertake several cleaning operations. Firstly, we drop all tweets that do not explicitly include a link to the URL. Second, users with fewer than 10 observed tweets over the time period were dropped. To further condense the dataset, we only include tweets which appeared at least 3 times in the dataset. As reposting and retweeting are important influence metrics on Twitter, this drops the tweets that likely did not receive as much attention. We also dropped domains that were only linked to by a single user and users that only linked to a single domain, as we hypothesized these pendulum nodes would be of limited use given the size of the graph. As a result of the expense of the webgraph attribute API, we chose to further restrict the Twitter data to only include tweets that mention at least one of the 43,758 domains for which we extracted attributes. This corresponds to keeping tweets that mention at least one of the 11,327 labeled domains for which we have attributes or tweets that mention one of the 11,327 and co-mention any of the 32,431 unlabeled domains. Following these cleaning operations, we are left with 555k users and 4.9M unique tweets which we observed tweeted or retweeted 91.2M times. This reduced dataset contains mentions of 2,475 reliability-labeled domains and 714 unlabeled back-linking domains.
3.4 Dredge Words
@@
With API constraints in mind, for the 100 websites with the lowest PC1 score rank in Lin et al. (2023). We extract the top-ranked $k\leq 1,000$ Google keyphrases from Ahrefs, for a total of 34,646 keyphrases. We then used WebSearcher Robertson and Wilson (2020) to query each of these keyphrases on Google and extracted the first 10 URLs returned in each SERP. We kept all queries for which the target unreliable domain was returned in the top 10 Google search results. This resulted in our set of 3,939 dredge words spanning 46 unreliable domains. We created a separate query to find mentions of these keyphrases in the 3.6B tweet covid dataset. This yielded 5.7 million tweets containing dredge words. Many of the most common mentions were explicit mentions of domain names or organization names (which can be identical to twitter handles—like “gatewaypundit”, “infowars”, and “nvic” all received over 10,000 mentions. To filter down the data, we use regular expressions to ensure each drudge word begins with a hashtag, starts a tweet, or is preceded by a white space. This retains 213 dredge words in 421k tweets that qualitatively contain less noise. Of these 421k tweets, only 9,788 (2%) explicitly linked to the unreliable domain associated with the dredge word.
@@
Some examples of dredge words include “psychic attack” (28 Twitter mentions), “akashic record” (12), “flu shot injury” (6), and “fallcabal” (5). While the majority of the dredge words are in English, other languages are present in the list, including Chinese, Hindi, and Arabic. We use this condensed twitter dataset and the set of SERP resultss they yielded in our dredge-word-based unreliable domain discovery process. While we were initially interested in paths from social media to search engines, we find dredge word SERPs surprisingly demonstrate a strong paths to social media. Youtube was by far the most commonly-returned domain (4,304 times). Wikipedia, Reddit, Quora, Twitter, Amazon, and Facebook were also among the most commonly-returned domains, in part due to the widespread commercialization of pseudo-scientific concepts. For a more in-depth analysis of dredge words see Appendix A.
4 Methods
In order to evaluate the impact of additional levels of context on our models, we construct graphs that capture increasingly-granular levels of context present in the data.
4.1 Single Mode Graphs
We define two single-mode graphs to evaluate the effectiveness of the models using only webgraph data and only Twitter data. Let $\mathcal{W}$ be the set of all labeled and unlabeled websites $\{w_{1},w_{2},\dots,w_{n}\}$ contained in the extracted webgraph data and define $\mathcal{U}$ as the set of all Twitter users $\{u_{1},u_{2},\dots,u_{n}\}$ that link to domains $\in\mathcal{W}$. For our webgraph data, we define $\mathcal{G}_{w}=(\mathcal{W},A,Z,Y)$ where $A$ is a binary undirected adjacency matrix where a $1$ or 00 denotes the presence or absence of a connection between website $i$ and website $j$. $Z$ is a node feature matrix, and $Y$ is the label assigned to each node ($1$ = reliable, 00 = unreliable, $-1$ = unlabeled). We define a corresponding graph for Twitter user-to-website connections $\mathcal{G}_{t}=(\mathcal{U},A,Z,Y)$ where nodes are users and websites $\in\mathcal{U}$ and the set of edges captured by $A$ denote observed edges between $\mathcal{U}$ and $\mathcal{W}$.

### v2

# Bridging Social Media and Search Engines: Dredge Words and the Detection of Unreliable Domains

@@ Abstract
Proactive content moderation requires platforms to rapidly and continuously evaluate the credibility of websites. Leveraging the direct and indirect paths users follow to unreliable websites, we develop a website credibility classification and discovery system that integrates both webgraph and large-scale social media contexts. We additionally introduce the concept of dredge words—terms or phrases for which unreliable domains rank highly on search engines—and provide the first exploration of their usage on social media. Our graph neural networks that combine webgraph and social media contexts generate to state-of-the-art results in website credibility classification and significantly improves the top-k identification of unreliable domains. Additionally, we release a novel dataset of dredge words, highlighting their strong connections to both social media and online commerce platforms.
Introduction
@@
On February 24, 2022, the day Russia began its invasion of Ukraine, Jacob Creech, a fringe Qanon conspiracy theorist, posted a series of unsubstantiated claims on Twitter. These tweets insinuated that the US had created COVID-19, and implied Russia’s invasion of Ukraine was actually being carried out to shut down US-funded “biolabs” in Ukraine and to prevent another global pandemic (League 2022). Hours after the tweet, the conspiratorial website InfoWars published an article promoting the tweet, crediting its author for uncovering an “ulterior motive theory” (League 2022). The tweet began circulating to other conspiratorial sites, and Google queries for “US biolabs” and “Ukraine Biolabs” spiked in the following days  [^1] https://trends.google.com/trends/explore?date=today“%205-y&geo=US&q=Ukraine“%20biolabs&hl=en. Conspiratorial sites were likely the only relevant search results until Snopes debunked the claim later that day (Evon 2022).
@@
In some ways, this is a success story of the current paradigm of combatting misinformation; fact-checkers responded rapidly to a quickly-spreading falsehood. In other ways, this highlights the challenges present in reactive misinformation interventions. Despite quick debunking by fact-checkers, the conspiracy was nonetheless echoed by established news sources like Fox News, and later amplified on the floor of the US Senate(League 2022). Researchers later found that a coordinated campaign of social media accounts boosted the narrative on Twitter (Alieva, Ng, and Carley 2022). As of 2024, despite the underlying claims having been debunked by many fact-checkers (Chappell and Yousef 2022; Parachini 2022), a search engine audit study found that Google, Bing, and Yandex all returned URLs on their first page of results that promote the “Ukraine Biolabs” conspiracy (Kuznetsova et al. 2024). The Ukraine Bioloab example highlights an important and understudied dimension of misinformation spread: the interaction between social media and search engines.
@@
An alternative to reactive fact-checking approaches are proactive algorithmic content moderation approaches. These can include modifying recommendation and ranking systems to decrease the reach and virality of unreliable information sources. In search engines, this can mean downranking articles from unreliable domains, and on social media, this can mean ranking posts containing unreliable domains lower in user newsfeeds. For proactive approaches to succeed, platforms need systems that can rapidly detect and discover unreliable domains. Both classification and detection are important as unreliable websites can, and do, deploy a number of tactics to avoid blacklists (Carragher, Williams, and Carley 2024a). In this work, we present a novel approach to the tasks of unreliable domain detection and discovery that leverages signals from both large-scale social media data and webgraph data.
@@
Directing users to search engines can be an effective tactic to spread misinformation as, since 2016, global surveys have consistently found that individuals report higher trust in search engines than traditional media (Barometer 2024; McDuling 2015). Thus by pointing users to search engines, either explicitly or implicitly, bad actors can create a false sense of content reliability. Little is understood about these pathways, and research has only recently begun exploring the ways social media users end up in data voids. The concepts of problematic queries (Golebiewski and Boyd 2019) and keyword signaling (Tripodi 2019) are highly related, but have only been explored in case studies. The recently-proposed concept of search directives provides a clear and observable path that could direct users to unreliable content on search engines. A search directive is content that is explicitly intended to prompt an online search, e.g., user $a$ directs user $b$ to ”look up Chemtrails on Google” (Robertson et al. 2023). In the context of data voids, search directives as a concept have two limitations: First, while search directives can direct users to unreliable content, they can also direct users to innocuous or helpful information, e.g., directing users to look up song lyrics or mathematical theorems. Second, as the search trends for “Ukraine Biolabs” demonstrate, users can be driven to unreliable content through search engines by querying phrases that appear without explicit search directives. To help fill these gaps, we propose the concept of dredge words—terms or keyphrases for which unreliable domains rank highly on search engines.
By attempting to explicitly incorporate each of the paths that users take to unreliable websites into GNNs, we seek to explore the dynamics that connect the spread of unreliable content on social media and search engines. We demonstrate that our relatively-simple curriculum-based heterogeneous graph model that leverage context from both webgraphs and social media data achieves SoTA results on the website credibility classification task. Further, our best model model does not incorporate any explicit text or semantic content from the webpages or from social media users, which makes this approach flexible, and easily-extendable to non-English contexts.
@@
Finally, we provide the first exploration of dredge words on social media. We incorporate a small set of dredge words into an unreliable domain discovery process in an attempt to mimic how social media users may transition from social platforms to the discovery of misinformation sources via search engines. Surprisingly, we find that dredge words frequently surface social media URLs in top Google Search Engine Result Page (SERP) positions—i.e., the websites returned when a user enters a query on Google Search. This suggests the existence of a bidirectional path that often leads back to social media. We show our heterogeneous model greatly outperforms competing systems in the top-k discovery of unlabeled unreliable websites. We publicly release the code, webgraph data collected for this project, a dataset of 3,939 dredge words for 46 unreliable news domains, and the resulting SERPs for each of the dredge word queries [^2] Artefacts will be publicly released following anonymous review period.. Our findings demonstrate that both direct and indirect paths to misinformation provide important signals that can be leveraged by researchers and platform designers working to mitigate the spread of misinformation across digital ecosystems.
Related Works
Social Media and Webgraphs
@@
Classifying individual texts or articles is a core problem addressed by research in misinformation detection. Such detection methods have typically relied on website content and social media data (Castelo et al. 2019; Chen and Freire 2020; Silva et al. 2021). Castelo et al. proposed a topic-agnostic detection system (Castelo et al. 2019) that identifies unreliable articles based on Linguistic Inquiry and Word Count features. Chen and Freire adopted the method for the task of unreliable domain discovery (Chen and Freire 2020). Their discovery system capitalized on user tendencies within the social graph, wherein a user who tweets a URL from a known unreliable source is likely to also tweet URLs from yet unknown sources. Similarly, Silva et al. combine website content and social context resulting in a misinformation page detection method that leverages heterogeneous data types (Silva et al. 2021), with a focus on early detection across a broad range of topics. The predictive power of webgraphs and social media data have been demonstrated separately in several contexts. Aswani et al. detect SEO manipulation by link-building sites by clustering on Pagerank score and Domain Authority (Aswani et al. 2021). For detecting news site bias, Aires et al. Patricia Aires, G. Nakamura, and F. Nakamura (2019) scrape cross-links from a list of prominent news sites from Media Bias Fact Check (Zandt 2022). Sehgal et al. explore a case where misinformation was spread through coordinated hyperlink and social media networks (Sehgal et al. 2021). Another recent case study demonstrated the manipulated webgraph linkages of unreliable pseudo-thinktanks (Williams and Carley 2023). Additionally, Zhang and Cabage show that social media and SEO promotion have different strengths; they find that social-sharing results in immediate but short-term boosts to traffic, while the benefits of link-building are slower to realize, but last longer (Zhang and Cabage 2017). This hints that a combination of both webgraph and social media data should be used in investigating misinformation sources. However, to our knowledge, this is the first work that attempts to combine the two contexts for unreliable domain classification or discovery.
Despite its importance and relevance, domain-level credibility detection is a relatively uncommon task in the literature—partially as a result of the lack of accepted labels. The most comprehensive label list as of 2024 are those published in Lin et al. (2023), which aggregates the scores of multiple different domain reliability rating lists. The vast majority of work in this space occurs at the article level (e.g., Nakov et al. (2023); Bianchi et al. (2024)). This class of approaches often take a sample of articles from each website and aggregate reliability ratings in some way. These approaches have several important limitations; namely, article-level approaches are restricted to languages with LLM support, model qualities depend heavily on what is sampled, many unreliable news sites often report true information with heavily partisan slants, and articles don’t necessarily reveal the goals of the publisher—e.g., knowing a media outlet has strong ties with a adversarial government can change how readers interpret the site’s content. Consequently, researchers have recently begun exploring content-agnostic domain classification approaches. In 2023, Yang and Menczer (2023) showed that LLMs could be used to evaluate website credibility. More recently, Carragher, Williams, and Carley (2024a) proposed a webgraph-based model for unreliable domain detection and discovery tasks. For discovery, the authors use backlinking domains in the webgraph in a snowball sampling approach. The authors demonstrate that GNN models trained on SEO attributes and webgraph data are effective for detection.
Data
We construct a heterogeneous graph with different relations based on the direct and indirect paths that users can take to unreliable websites. To capture direct paths to target news sites, we collect the 10 domains which most frequently link to each labeled news site (e.g., website $i$ has linked 1M times to website $j$). We further extract social media mentions of each target news site (e.g., user $k$ posts a hyperlink to website $j$). We consider both of these to be direct paths, as in both these cases, users would simply click a link to access website $j$. Finally we consider an indirect path through dredge words. Dredge words connect users to websites through Google SERPs. When browsing social media, users can encounter words or concepts with which they’re unfamiliar and that exist in data voids—for example, a user might see a Twitter post about “indigo children” and search a variation of that query on Google. The results surface numerous pseudoscientific websites and advertisements (see Appendix A). We provide a summary visualization of this heterogeneous network in Figure 1.
Figure 1: A summary of the heterogeneous graph construction process. Solid lines denote direct paths (a user clicks a hyperlink), and dashed lines denote indirect paths (a user sees a post and then queries a subset of that post on a search engine).
Webgraph and features
We use the domain credibility labels published by Lin et al. (2023). The authors of the work align the domain reliability ratings of all domains ranked by 6 expert groups and run imputation followed by principal component analysis to generate aggregate ratings for 11,520 domains. We binarize the news domain rankings as ”reliable” and ”unreliable” using a threshhold of 0.5162, corresponding to the bottom two quintiles of the data. Following the methodology of Carragher, Williams, and Carley (2024a), We use the SEO toolkit service Ahrefs333ahrefs.com to extract the 10 domains which link to each of the 11,520 target domains at the highest volume (the highest-volume back-linking domains). Ahrefs feature values have been found to correlate strongly with ground-truth pagerank calculations and competing traffic estimate systems (Carragher, Williams, and Carley 2024b). In total, we extracted 43,758 domains, each with 23 domain-level attributes. Additional discussion of labels, collection details, and EDA of the webgraph can be found in Appendix A. Each edge in this dataset denotes a direct path, as a user on the source site could click a hyperlink to move to the reliability-labeled target site. We were unable to extract features for 193 domains—the domains largely appeared dead or inactive, so we excluded them from the data. This left us with 11,327 labeled domains.
Twitter Data
@@
For social media context, we use a Twitter dataset constructed by querying COVID-related keywords444coronavirus, Wuhan virus, Wuhanvirus, 2019nCoV, NCoV, NCoV2019, covid-19, covid19, covid 19 via Twitter’s streaming API between January 29, 2020 and June 26, 2022. Due to server issues and API limitations, 121 days over the time period have partial or missing data. However, these gaps are spread relatively evenly over the time period, and so the data still provide strong coverage. Our final Twitter dataset contained 3.6 billion extracted tweets.
@@
From this set of 3.6 billion tweets, we extract all Tweets that mention or link to one of our 11,327 websites, resulting in a dataset of 320M tweets that link to over 840k unique domains. Of the original 11,327 labeled domains, only 5,504 appeared within the Twitter data. To reduce the noise in the dataset we undertake several cleaning and filtration operations, which, due to length constraints, we detail in Appendix A. Following these cleaning operations, we are left with 555k users and 4.9M unique tweets which we observed tweeted or retweeted 91.2M times. This reduced dataset contains mentions of 2,475 reliability-labeled domains and 714 unlabeled back-linking domains.
Dredge Words
@@
For the 100 websites with the respective lowest and highest reliability rankings, we extract the top-ranked $k\leq$ 1,000 Google keyphrases from Ahrefs, for a total of 34,646 keyphrases. We then used WebSearcher (Robertson and Wilson 2020) to query each of these keyphrases on Google and extracted the first 10 URLs returned in each SERP. We kept all queries for which the target unreliable domain was returned in the top 10 Google search results. As many of the domains for which we collected dredge words did not have any queries for which they ranked in the top 10 on Google, this resulted in a set of 3,939 dredge words spanning 46 unreliable domains. We created a separate query to find mentions of these keyphrases in the 3.6B tweet covid dataset. This yielded 5.7 million tweets containing dredge words. Many of the most common mentions were explicit mentions of domain names or organization names (which can be identical to twitter handles—like “gatewaypundit”, “infowars”, and “nvic” all received over 10,000 mentions. To filter down the data, we use regular expressions to ensure each drudge word begins with a hashtag, starts a tweet, or is preceded by a white space. This retains 213 dredge words in 421k tweets that qualitatively contain less noise. Of these 421k tweets, only 9,788 (2%) explicitly linked to the unreliable domain associated with the dredge word.
@@
This data pipeline attempts to mirror the way a user might access this information “in the wild”. A user might encounter a Twitter post that contains a reference to “silent assassination through amplified neurons”, or “Project S.A.T.A.N.”, search the query on Google, and encounter unreliable results like those in Figure 2. Some additional examples of dredge words include “psychic attack” (28 Twitter mentions), “akashic record” (12), “flu shot injury” (6), and “fallcabal” (5). However, we note that our current collection pipeline is almost certainly underestimating their actual usage as 1) on Twitter, we only capture dredge word usage mentioned alongside COVID keywords, and 2) we extracted dredge words in 2024, 2-4 years after tweets were posted. While the majority of the dredge words are in English, other languages are present in the list, including Chinese, Hindi, and Arabic. We use this condensed twitter dataset and the set of SERP results they yielded in our dredge-word-based unreliable domain discovery process. While we were initially interested in paths from social media to search engines, we find dredge word SERPs surprisingly demonstrate a strong paths to social media. Youtube was by far the most commonly-returned domain (4,304 times). Wikipedia, Reddit, Quora, Twitter, Amazon, and Facebook were also among the most commonly-returned domains, in part due to the widespread commercialization of pseudo-scientific concepts.
Case Study
To provide a more concrete illustration of the tasks and our motivation, we’ll consider some of the ways that users might end up on the unreliable website, Gaia [^5] https://mediabiasfactcheck.com/gaia/. This website often promotes pseudoscience, vaccine misinformation, and various conspiracy theories. We can imagine two users taking direct paths and a third following an indirect path. The first imaginary user follows paranormal accounts or pages on social media, and sees an account post a link to an article about haunted houses on Gaia. The second individual is interested in UFOs and listens to the “Coast to Coast AM” [^6] https://en.wikipedia.org/wiki/Coast˙to˙Coast˙AM radio talk show—if that individual visited the website for the talk show, they would find 29.4K hyperlinks that would lead to Gaia. In a third instance, we can imagine a user interested in astrology seeing the post in Figure 3.
Figure 3: A tweet about “Indigo Children”.
The user, who is encountering new information, might then be curious and compelled to search “indigo children meaning” or just “indigo children” on Google—both phrases are dredge words that surface Gaia in the top 10 results. Interestingly, after manually searching these “indigo children meaning” in June 2024, we found that Google’s definition snippet—a summarization box on the top of some SERPs with the text “From sources across the web”—expresses pseudo-scientific concepts. The box states that “Indigo Children are kids with indigo-colored auras”. This snippet was followed by YouTube videos promoting Indigo Children pseudoscience and a Reddit post from the /r/Psychic subreddit. Another user who may have searched only “indigo children” would have seen a snippet from Wikipedia which correctly calls indigo children pseudoscientific. However, adjacent to the snippet, the user would have seen four books for sale about indigo children. After that, the second highest-ranking return was an article from an unreliable domain identifying the “13 signs you’re an indigo child”.
This case study also highlights the importance of conceptualizing misinformation consumption as interconnected; none of our hypothetical users were actively seeking anti-vaccine content, and yet in pursuing their relatively-innocuous interests, all are exposed to medical misinformation. While all of these paths lead to Gaia, the indirect routes may be the most effective at swaying users, as discovering Gaia through a trusted intermediary (in this case, Google) may confer additional trust to the source (Robertson et al. 2023). The questions we are trying to answer in this work, with respect to this case study, are 1) “how does integrating each of these path contexts into a model impact our ability to identify Gaia as unreliable?” and 2) “What websites that link to Gaia are also promoting misinformation”? Even in this case study, neither tasks are trivial; detecting unreliable websites is challenging, and while many of the websites that link to Gaia are unreliable, others are generic SEO sites. Additionally, there are reliable and .edu sites that link to Gaia in fact-checking articles, which further complicates the task.
Methods
In order to evaluate the impact of additional levels of context on our models, we construct graphs that capture increasingly granular levels of context present in the data. Let $\mathcal{W}$ be the set of all labeled and unlabeled websites $\{w_{1},w_{2},\dots,w_{n}\}$ contained in the extracted webgraph data. We define $\mathcal{W}_{target}$ as the subset of $\mathcal{W}$ for which we have reliability labels—i.e., the 11,327 reliability-labeled news domains; we define $\mathcal{W}_{source}$ as the 10 websites that most link to each domain. We note that $\mathcal{W}_{source}$ contains some websites in $\mathcal{W}_{target}$. Next define $\mathcal{U}$ as the set of all Twitter users $\{u_{1},u_{2},\dots,u_{n}\}$ that link to domains $\in\mathcal{W}$. Finally, let dredge words $\mathcal{Q}$ be the set of keyphrases $\{q_{1},q_{2},\dots,q_{n}\}$ connected to SERPs containing URLs $\in\mathcal{W}$.
We define several preliminaries. A Graph Union operation ($G=G_{1}\cup G_{2}$), between graphs $G_{1}=(V_{1},E_{1})$ and $G_{2}=(V_{2},E_{2})$ is defined as $G=(V,E)$ where $V=V_{1}\cup V_{2}$ and $E=E_{1}\cup E_{2}$. A heterogeneous graph is an extension of a homogeneous graph, $\mathcal{G}=(V,E)$, where $V$ and $E$ are associated with a node type mapping function $\Psi:V\rightarrow A$ and an edge type mapping function $\Phi:E\rightarrow\phi$. In our setting, the set of node types are $A=\{\mathcal{W},\mathcal{U},\mathcal{Q}\}$ and the set of edge types are $\Phi=\{\phi_{1}:domain-to-domain,\phi_{2}:user-to-domain,\phi_{3}:user-to-dredgeword,\phi_{4}:dredgeword-to-domain\}$. This setup is visualized in Figure 1.
Homogeneous Graphs
We consider three homogeneous baselines—where all nodes are treated as part of the same class—to justify the added complexity of using heterogeneous networks. We define two homogeneous graphs: $\mathcal{H}_{domains}$ contains domain-to-domain relationships, i.e., $\mathcal{W}_{source}$ to $\mathcal{W}_{target}$. $\mathcal{H}_{users}$ contains relationships between Twitter users and the websites they mentioned $\in\mathcal{W}$. Finally, we define $\mathcal{H}_{domains+users}$ = $\mathcal{H}_{domains}\cup\mathcal{H}_{users}$. Clearly, these latter two graphs contain different sets of nodes and relations, but we treat them as equivalent in these experiments as a simple baseline.